# Back from the dead: AT2019aalc as a candidate repeating TDE in an AGN

### v2

@@ 5.5 SED fitting
Following Reusch et al. (2022), we fitted two blackbody models in order to describe the SED of the transient using lmfit. These fitted curves to the optical/UV (’blue’ blackbody) and IR datapoints (’red’ blackbody) are shown in Fig. 15, together with a combined double blackbody fit. The red blackbody fit was introduced to account for the IR dust echo. These curves fit the SED around the optical peak of the investigated flare.
The continuum emission of TDEs is well described by a thermal blackbody model (Gezari 2021). The case of AT2019aalc appears to be more complex. The UV emission has a contribution from the Bowen Fluorescence mechanism and possibly also from Comptonization explaining the poor fit towards the blue. The excess seen in the WISE bands is possibly due to the non-simultaneous observation with respect to the optical/UV and WIRC observations as the IR light curve is peaked well after the optical one. The excess in the -band possibly has a contribution from the early appearance of some Fe lines in this range but such a significant excess cannot be fully explained with the coronal lines only.
The inferred blackbody temperature and radius values estimated for the ’blue’ blackbody fit are K and cm, respectively. The blackbody temperature is in the lower range of that is found for optical selected TDEs, however, an order of magnitude lower than the typical values of the X-ray detected TDEs studied by van Velzen et al. (2020) and Thomsen et al. (2022). On the other hand, the radius is smaller than calculated for optical selected TDEs but is more consistent with the X-ray detected ones. The ’red’ blackbody fit resulted in K and cm.
@@ 5.6 What type of transient is AT2019aalc?
AT2019aalc shares similarities with TDEs such as the soft X-ray spectrum, the dust echo and the optical flare(s) rising with constant color. The UV bright nature of the transient with a blue optical/UV peak color of NUV is another TDE-like feature and not typical for AGN and SNe (Gezari 2021). Bowen lines appear in the spectra of severeal TDEs even if these are typically weaker than in our case. The most remarkable difference is, however, the fading of the optical emission, which is slower than for any TDE identified so far. The reddening in color when the main optical flares are declining is also atypical for TDEs.
@@
The unobscured AGN-like optical spectra, the presence of persistent and strong Bowen lines (O iii and N iii ), the extremely luminous UV flare, on yearly timescales decaying optical emission and the optical re-brightening altogether suggest a Bowen Fluorescence Flare classification of AT2019aalc. Furthermore, we classify AT2019aalc as an Extreme Coronal Line Emitter because of the detection of several strong high-ionization coronal lines.
@@
The mechanism that powers the BFFs is still unclear. TDEs in AGN might play a role. In a single star case, the slowly decaying optical/UV emission might be explained by the dissipation of energy due to the continuous interaction between the stellar debris stream and the pre-existing accretion disk (similarly to the case of the TDE-AGN PS16dtm, Blanchard et al. 2017, and ASASSN-14ko, Huang et al. 2023) or the AGN’s radiation field. According to TDE-AGN simulations, stellar passages can induce substantial perturbations in the inner disk and leave it in a very perturbed state even for years (Chan et al. 2019). The debris-disk interaction results in enhanced accretion and multiple shocks (Chan et al. 2019), notably leading to X-ray emission (Ryu et al. 2024). Consistently within this picture, the soft X-ray flares of AT2019aalc suggest that the emission is linked to the accretion disk. These X-ray flares lead the optical/UV bumps by weeks, which are detected during the decay of the main flare. The soft X-ray emission is likely reprocessed in optical/UV by the dense environment explaining the optical/UV flares and the emergence of the Bowen fluorescence and coronal lines. The TDE-AGN scenario provides an explanation for one slowly decaying flare with recurring outbursts, however, AT2019aalc exhibits two distinct optical flares. TDE-AGN simulations also revealed that in a case of an initially partial disruption, a surviving core would generate similar heating of the inner disk to the first pericenter passage (Chan et al. 2019; Ryu et al. 2024). Notably, the second flare of AT2019aalc is more luminous than the first one, similarly to the cases of the partial TDE AT2020vdq (Somalwar et al. 2023) and the BFF F01004-2237. Numerical simulations demonstrated that the even more luminous second flare of a partial TDE might be explained in the following way: tidal energy injection leads to oscillations in the remnant star, and if the thermal energy cannot dissipate within one orbital period, the star expands, making it more susceptible to further tidal disruption and potentially resulting in brighter flares (Zhong 2024; Liu et al. 2024). Liu et al. (2024) claimed that these considerations provide a plausible explanation for the brighter second flare of the partial TDE AT2020vdq. A partial TDE scenario might explain the second optical flare of the other two BFFs (AT2017bgt and AT2021loi) with double-peaked optical flares as well.
@@
Although the slowly decaying optical emission and the re-brightening seem to be common features of BFFs, the long-term light curves still show differences as well. The re-brightening episodes started at different timescales for the individual sources, and more BFFs need to be classified to better understand their unique optical light curves. Nevertheless, (partial) TDEs occurring in AGN can be a possible origin of BFFs. This would explain the TDE-like properties of these sources. The most remarkable differences, such as the slowly decaying optical emission and the continuum emission not being well described by blackbody models, may arise due to the more complex environment in which the TDE occurs, compared to the simpler conditions in quiescent galaxies (”TDEs in vacuum”). TDE-AGN simulations showed that the emergent spectrum of such a system may not be purely thermal as in regular TDEs (Chan et al. 2019). Furthermore, a possible connection between the high-ionization coronal lines (which are most probably related to TDE activity, Wang et al. 2012) and the BFFs (see Fig.13) is consistent with this picture. A known candidate partial TDE in an AGN is the case of the active galaxy IC 3599. This low-luminosity AGN exhibits periodic soft X-ray flares interpreted with a partially disrupted star, which results in reoccurring outbursts after each passage of the surviving core around the central black hole (Campana et al. 2015). Notably, IC 3599 has more prominent high-ionization coronal lines than most AGN (similarly to the coronal-line detected transients, see Fig. 13). In addition, IC 3599 is the only known AGN with fading coronal lines (Frederick et al. 2019).
@@ 5.7 Radio properties of BFFs
As we studied AT2019aalc in radio in detail, we compare the radio properties with that of the BFFs. Based on the radio observations explained in Makrygianni et al. (2023), AT2021loi has an inverted radio spectrum between and GHz while the source was not detected at GHz. Although no radio spectrum is available for AT2017bgt, we investigated the VLASS QL Images of this BFF to study its radio flux variability. We found a radio source at the optical position of AT2017bgt with flux densities of mJy (on 2019-05-21), mJy (on 2021-10-19) and mJy (on 2024-06-11) at a central frequency of GHz. The pre-flare system was detected in the FIRST (at GHz) in July 1998 with a flux density of mJy. The BFF AT2020afhd was detected with the VLA at GHz with mJy (Christy et al. 2024). If BFFs are explainable with TDE-AGN, the synchrotron emission is likely originating from outflow activity due to the enhanced accretion or, alternatively, from interaction between the unbound debris and the surrounding gas predicted by TDE-AGN simulations published in Chan et al. (2019). Radio observations of any newly discovered BFF will be essential to better understand these transients.
@@ 5.8 BFFs and high-energy neutrinos
AT2019aalc was associated with the high-energy neutrino event IC191119A (most probable neutrino energy: TeV) detected by the IceCube Observatory (IceCube Collaboration 2019; van Velzen et al. 2024). Motivated by the classification of AT2019aalc suggested by its multi-wavelength properties (Subsect. 5.6), we searched for spatial coincidences between the earlier classified BFFs and the neutrino alerts from the IceCube Event Catalog of Alert Tracks (ICECAT-1: Abbasi et al. 2023). We found that the BFF AT2021loi is located within the 90% rectangular uncertainty contour of the high-energy neutrino event IC230511 (most probable neutrino energy: TeV) detected in May (IceCube Collaboration 2023). In the full likelihood contours published in ICECAT-1, the source is just outside the 90% contour. We present the source positions together with the neutrino best-fit positions and both contours in Fig 16. We also report a coincidence between the ’prototype’ BFF AT2017bgt and the high-energy neutrino event IC-200410A. This event is, however, a poorly reconstructed cascade event, therefore we do not discuss it further here.
The neutrino associated with AT2021loi was detected days and days after the first and second optical peaks, respectively. Notably, all three nuclear transients (including AT2019aalc) associated with neutrinos exhibited a similar pattern, with neutrino detection occurring days after the optical peaks. This might be explained with delayed outflow activity (van Velzen et al. 2024). In the case of AT2021loi, the inverted radio spectrum (Makrygianni et al. 2023) is compatible with a newly ejected outflow component, similarly to the case of AT2019aalc. Interactions between the out-flowing material and clouds in the surrounding of the SMBH (Wu et al. 2022) or between ultra-high energy protons (accelerated in a jet or outflow) and infrared photons from the dust echo (Winter & Lunardini 2023) were proposed to explain the observed neutrino emission of the nuclear transients. These findings indicate that outflows play an important role in the neutrino production. A case study for AT2021loi would help us to further constrain the neutrino production models of nuclear transient sources.
If the connection between the high-ionization coronal line detection and the unusually large dust echos of nuclear transients is real, this can be used to reveal more sources with strong dust echo and therefore more candidate neutrino sources. Notably, both AT2019aalc and AT2021loi are characterized by significant dust echos (van Velzen et al. 2024; Hinkle 2024: see also Subsect. 4.2) and unusually strong high-ionization coronal lines (Makrygianni et al. 2023: see also Subsect. 5.1).
@@ 6 Conclusions
We conducted a multi-wavelength campaign during the re-brightening of the nuclear transient AT2019aalc associated with a high-energy neutrino event. We found that the second optical flare is more luminous and decays slower than the first one. Both flares evolve rapidly and with constant color. The UV emission peaks at extremely high luminosities (at erg/s). The re-brightening is accompanied by an IR flare, similarly to the first flare when it was explained with dust echo emission. An X-ray flare peaked before the optical/UV emission, preceding two subsequent flares that were nearly an order of magnitude brighter and predominantly blackbody in nature. The X-ray flares are very soft which explains the unusually strong He ii lines. In the optical spectra, persistent Bowen fluorescence lines were detected over a year after the second flare. With VLBI we clearly detect a non-thermal radio source which favors an AGN-like origin. The radio flux monitoring revealed a long-term flare and an inverted spectrum above GHz, which is likely explained by a newly ejected, self synchrotron-absorbed outflow component. Altogether, our results suggest a Bowen Fluorescence Flare classification of AT2019aalc. We also classify AT2019aalc as an Extreme Coronal Line Emitter.
The two distinct optical flares might be attributed to the partial disruption of a single star within an AGN environment. The stellar debris-disk interaction in such a system could result in complex light curve variations e.g., the plateau observed in optical/UV (Chan et al. 2020). A TDE-related scenario provides an explanation not only for the suddenly enhanced activity of the nucleus but also for the high metallicity of the disk gas which was notably suggested by TDE-AGN simulations (Ryu et al. 2024). Future observations will help us to better understand the nature of AT2019aalc: a repeating TDE would lead to a roughly periodic flaring, while some other AGN activity would lead to erratic future flares.
Motivated by the high-energy neutrino associated with AT2019aalc, we cross-matched the ICECAT-1 with the positions of the known 4 BFFs. We found that AT2021loi is coincident with the high-energy neutrino event IC230511 confirming the flaring AGN-neutrino connection studied by van Velzen et al. (2024). Apart from the known multi-wavelength properties of BFFs, AT2021loi shares characteristics with AT2019aalc such as late-time IR flare (due to dust echo), high-ionization coronal lines in its spectra (and also -band excess), radio detection (and inverted radio spectrum) and re-brightening episode in optical. In both cases the AGN are classified as broad-line AGN in-prior the transients occurred.

### v3

@@ 5.5 SED fitting
Following Tywin, we fitted two blackbody models in order to describe the SED of the transient using lmfit. These curves fitted to the optical/UV (blue blackbody) and IR datapoints (red blackbody) are shown in Fig. 15, together with a combined double blackbody fit. The red blackbody fit was introduced in order to account for the IR dust echo. These curves fit the SED around the optical peak of the investigated flare.
The continuum emission of TDEs is described well by a thermal blackbody model (2021ARA&A..59...21G). The case of AT2019aalc appears to be more complex. The UV emission has a contribution from the Bowen Fluorescence mechanism and possibly also from Comptonization, explaining the poor fit towards the blue. The excess seen in the WISE bands is possibly due to the non-simultaneous observation with respect to the optical/UV and WIRC observations, as the IR light curve is peaked well after the optical one. The excess in the -band possibly has a contribution from the early appearance of some Fe lines in this range, but such a significant excess cannot be fully explained with the coronal lines only. Instead, a variable H component, given its much larger equivalent width, or reprocessing, is more likely to significantly contribute to the observed excess and warrants consideration.
The inferred blackbody temperature and radius values estimated for the blue blackbody fit are K and cm, respectively. The blackbody temperature is in the lower range of that is found for optical selected TDEs, however, an order of magnitude lower than the typical values of the X-ray detected TDEs studied by 2020SSRv..216..124V and 2022ApJ...937L..28T. On the other hand, the radius is smaller than those calculated for optically selected TDEs, however is more consistent with the X-ray detected ones. The red blackbody fit resulted in K and cm.
@@ 5.6 What type of transient is AT2019aalc?
AT2019aalc shares similarities with TDEs including the soft X-ray spectrum, the dust echo and the optical flare(s) rising with constant color. The UV bright nature of the transient with a blue optical/UV peak color of NUV is another TDE-like feature and not typical for AGN and SNe (2021ARA&A..59...21G). Bowen lines appear in the spectra of several TDEs even if these are typically weaker than in our case. The most remarkable difference is, however, the fading of the optical emission, which is slower than for any TDE identified so far. The reddening in color when the main optical flares are declining is also atypical for TDEs. However, in a more complex AGN environment, compared to the simpler case of standard TDEs in quiescent galaxies, reprocessing may already be significant at early times, implying that the observed reddening does not necessarily reflect genuine cooling.
@@
The unobscured AGN-like optical spectra, the presence of persistent and strong Bowen lines (O iii and N iii ), the extremely luminous UV flare, optical emission decaying on yearly timescales, and the optical re-brightening altogether suggest a Bowen Fluorescence Flare classification of AT2019aalc. Furthermore, we classify AT2019aalc as an Extreme Coronal Line Emitter because of the detection of several strong high-ionization coronal lines.
@@
The mechanism that powers the BFFs is still unclear. TDEs in AGN might play a role. In a single star case, the slowly decaying optical/UV emission might be explained by the dissipation of energy due to the continuous interaction between the stellar debris stream and the pre-existing accretion disk (similarly to the case of the TDE-AGN PS16dtm, 2017ApJ...843..106B, and ASASSN-14ko, 2023ApJ...956L..46H) or the AGN’s radiation field. According to TDE-AGN simulations, stellar passages can induce substantial perturbations in the inner disk and leave it in a very perturbed state even for years (tdes_in_agn). The debris-disk interaction results in enhanced accretion and multiple shocks (tdes_in_agn), notably leading to X-ray emission (2024MNRAS.527.8103R). Consistently within this picture, the soft X-ray flares of AT2019aalc suggest that the emission is linked to the accretion disk. These X-ray flares lead the optical/UV bumps by weeks, which are detected during the decay of the main flare. The soft X-ray emission is likely reprocessed to optical/UV by the dense environment, explaining the optical/UV flares and the emergence of the Bowen fluorescence and coronal lines. The TDE-AGN scenario provides an explanation for one slowly decaying flare with recurring outbursts, however, AT2019aalc exhibits two distinct optical flares. TDE-AGN simulations also revealed that in a case of an initially partial disruption, a surviving core would generate similar heating of the inner disk to the first pericenter passage (tdes_in_agn; 2024MNRAS.527.8103R). Notably, the second flare of AT2019aalc is more luminous than the first one, similarly to the cases of the partial TDE AT2020vdq (2020vdq) and the BFF F01004-2237. Numerical simulations demonstrated that the even more luminous second flare of a partial TDE might be explained in the following way: tidal energy injection leads to oscillations in the remnant star, and if the thermal energy cannot dissipate within one orbital period, the star expands, making it more susceptible to further tidal disruption and potentially resulting in brighter flares (vdq2; partial_models). vdq2 claimed that these considerations provide a plausible explanation for the brighter second flare of the partial TDE AT2020vdq. A partial TDE scenario might explain the second optical flare of the other two BFFs (AT2021loi and F01004-223) with double-peaked optical flares as well.
@@
A TDE-in-AGN scenario would explain the TDE-like properties of these sources. The most remarkable differences, such as the slowly decaying optical emission and the continuum emission not being well described by blackbody models, may arise due to the more complex environment in which the TDE occurs, compared to the simpler conditions in quiescent galaxies (”TDEs in vacuum”). TDE-AGN simulations showed that the emergent spectrum of such a system may not be purely thermal as in regular TDEs (tdes_in_agn). Furthermore, a possible connection between the high-ionization coronal lines (which are most probably related to TDE activity, 2012ApJ...749..115W) and the BFFs (see Fig. 13) is consistent with this picture. A known candidate partial TDE in an AGN is the case of the active galaxy IC 3599. This low-luminosity AGN exhibits periodic soft X-ray flares interpreted as being due to interaction with a partially disrupted star, which results in reoccurring outbursts after each passage of the surviving core around the central black hole (2015A&A...581A..17C). Notably, IC 3599 has more prominent high-ionization coronal lines than most AGN (similarly to the coronal-line detected transients, see Fig. 13). In addition, IC 3599 is the only known AGN with fading coronal lines (2019ApJ...883...31F) a behavior also observed in the TDE AT2019qiz (2023MNRAS.525.1568S), the BFF F01004-2237 (f01) and most of the ECLEs (2013ApJ...774...46Y). Another feature AT2019aalc shares with other candidate pTDEs is the Heii + Niii complex, observed in the optical spectra of all five optically selected pTDE candidates (see Fig. 13. in 2024arXiv241215326H).
Interestingly, other BFFs also showed weaker pre-transient AGN activity than persistently accreting AGN (Trakhtenbrot; AT2021loi). It suggests that TDEs might have reactivated these AGN. However, between the two main optical flares, AT2019aalc exhibited negative fluxes, i.e., it was fainter than the baseline (see Fig. 2). This behavior suggests a retrograde orbit of the disrupted star relative to the accretion disk, as such encounters gain angular momentum and lead to a subsequent lower AGN state (2022MNRAS.514.4102M). In contrast, a prograde orbit adds angular momentum to the disk, resulting in a higher AGN state, which may explain the light curve of OGLE17aaj, where the source persisted in a high and stable state for years after the peak (2019A&A...622L...2G), unlike AT2019aalc.
@@ 5.7 Radio properties of BFFs
As we studied AT2019aalc in radio in detail, we compare the radio properties with that of the BFFs. Based on the radio observations explained in AT2021loi, AT2021loi has an inverted radio spectrum between and GHz while the source was not detected at GHz. Although no radio spectrum is available for AT2017bgt, we investigated the VLASS QL Images of this BFF to study its radio flux variability. We found a radio source at the optical position of AT2017bgt with flux densities of mJy (on 2019-05-21), mJy (on 2021-10-19) and mJy (on 2024-06-11) at a central frequency of GHz. The pre-flare system was detected in the FIRST (at GHz) in July 1998 with a flux density of mJy. F01004-2237 was detected with VLBI at Ghz, however, remained undetected at GHz which indicates an inverted spectrum (2024ApJ...970....5H). The BFF AT2020afhd was detected with the VLA at GHz with mJy (2024TNSAN..56....1C). If BFFs are explainable with TDE-AGN, the synchrotron emission is likely originating from outflow activity due to the enhanced accretion or, alternatively, from interaction between the unbound debris and the surrounding gas predicted by TDE-AGN simulations (tdes_in_agn). Radio observations of any newly discovered BFF will be essential to better understand these transients.
@@ 5.8 BFFs and high-energy neutrinos
AT2019aalc was associated with the high-energy neutrino event IC191119A (most probable neutrino energy: TeV) detected by the IceCube Observatory (2019GCN.26258....1I; Lancel). Motivated by the classification of AT2019aalc suggested by its multi-wavelength properties (Subsect. 5.6), we searched for spatial coincidences between the earlier-classified BFFs and the neutrino alerts from the IceCube Event Catalog of Alert Tracks (ICECAT-1: 2023ApJS..269...25A). We found that the BFF AT2021loi is located within the 90% rectangular uncertainty contour of the high-energy neutrino event IC230511 (most probable neutrino energy: TeV) detected in May (2023GCN.33773....1I). In the full likelihood contours published in ICECAT-1, the source is just outside the 90% contour. We present the source positions together with the neutrino best-fit positions and both contours in Fig 16. We also report a coincidence between the prototype BFF AT2017bgt and the high-energy neutrino event IC-200410A. This event is, however, a poorly reconstructed cascade event, therefore we do not discuss it further here.
The neutrino associated with AT2021loi was detected days and days after the first and second optical peaks, respectively. Notably, all three nuclear transients (including AT2019aalc) associated with neutrinos exhibited a similar pattern, with neutrino detection occurring days after the optical peaks. This might be explained with delayed outflow activity (Lancel). In the case of AT2021loi, the inverted radio spectrum (AT2021loi) is compatible with a newly ejected outflow component, similarly to the case of AT2019aalc. Interactions between the out-flowing material and clouds in the surrounding of the SMBH (2022MNRAS.514.4406W) or between ultra-high-energy protons (accelerated in a jet or outflow) and infrared photons from the dust echo (2023ApJ...948...42W) were proposed to explain the observed neutrino emission of the three nuclear transients. These findings indicate that outflows play an important role in the neutrino production. A case study for AT2021loi would help us to further constrain the neutrino production models of nuclear transient sources.echos
If the connection between the high-ionization coronal line detection and the unusually large dust echos of nuclear transients is real, this can be used to reveal more sources with strong dust echoes and therefore more candidate neutrino sources. Notably, both AT2019aalc and AT2021loi are characterized by significant dust echoes (Lancel; 2024MNRAS.531.2603H: see also Subsect. 4.2) and unusually strong high-ionization coronal lines (AT2021loi: see also Subsect. 5.1).
@@ 6 Conclusions
We conducted a multi-wavelength campaign during the re-brightening of the nuclear transient AT2019aalc, which have been associated with a high-energy neutrino event. We found that the second optical flare was more luminous and decayed slower than the first one. Both flares rose rapidly and with constant color. The UV emission peaked at extremely high luminosities (at erg/s). The re-brightening was accompanied by an IR flare, similarly to the first flare, when it was explained with a dust echo emission. An X-ray flare peaked before the optical/UV emission, preceding two subsequent flares that were nearly an order of magnitude brighter and predominantly blackbody in nature. The X-ray flares are very soft, which explains the unusually strong He ii lines. In the optical spectra, persistent Bowen fluorescence lines were detected over a year after the second flare. With VLBI, we clearly detect a non-thermal radio source, which favors an AGN-like origin. The radio flux monitoring revealed a long-term flare and an inverted spectrum above GHz, which is likely explained by a newly ejected, self synchrotron-absorbed outflow component. Altogether, our results suggest a Bowen Fluorescence Flare classification of AT2019aalc. We also classify AT2019aalc as an Extreme Coronal Line Emitter.
The two distinct optical flares might be attributed to the partial disruption of a single star within an AGN environment. The stellar debris-disk interaction in such a system could result in complex light curve variations, e.g., the plateau observed in optical/UV (tdes_in_agn_2). A TDE-related scenario provides an explanation for the suddenly enhanced activity of the nucleus. Future observations will help us to better understand the nature of AT2019aalc: a repeating TDE would lead to a roughly periodic flaring, while some other AGN activity would lead to erratic future flares.
Motivated by the high-energy neutrino associated with AT2019aalc, we cross-matched the ICECAT-1 with the positions of the four known BFFs. We found that AT2021loi is coincident with the high-energy neutrino event IC230511, confirming the flaring AGN-neutrino connection studied by Lancel. Apart from the known multi-wavelength properties of BFFs, AT2021loi shares characteristics with AT2019aalc such as late-time IR flare (due to dust echo), high-ionization coronal lines in its spectra (and also -band excess), radio detection (and inverted radio spectrum) and a re-brightening episode in optical. In both cases the AGN are classified as broad-line AGN prior to when the transients occurred.